Please answer the following showing your steps on how you got the answer.
Use the explicit formula a[n]=a[1]+(n-1)d to find the 150th term of the sequence below.
19,25,31,37,43, ...

Answer by stanbon(75887)

Use the explicit formula a[n]=a[1]+(n-1)d to find the 150th term of the sequence below.
19,25,31,37,43, ...
-----------------------------------
a(1) = 19
d is the difference between any two terms: 25-19 = 6
a(n) is the nth term
--------------------------------
Ans: a(150) = 19 + (149)6 = 19 + 894 = 913
===========================================
Cheers,
Stan H.